OpenStudy (anonymous):

If x,y > 0, log_y (x) + log_x (y) = 10/3 and xy = 144, then (x + y)/2 = Answer choices: available upon request. (How do you get the answer?)

OpenStudy (anonymous):

i will byte u :P

OpenStudy (anonymous):

\[\left( \log_yx\right)^2 + 1 = \frac{10}3\log_yx\]Complete squares to get the value of \(\log_yx\).

OpenStudy (anonymous):

ops bite :p

OpenStudy (anonymous):

@Ishaan94 is right :) Solve using quadratic formula the equation he wrote

OpenStudy (anonymous):

log_y (x) = log_x (x)/log_x(y) so: log_x (x)/log_x(y) +log_x(y) = 10/3 log_x(x/y *y) =10/3 x=10/3 rest should be easy

OpenStudy (anonymous):

@myko , log_x (x)=1

OpenStudy (anonymous):

:) good point

OpenStudy (anonymous):

@myko , if you substitute that, you will get same equation as @Ishaan94 got

OpenStudy (anonymous):

ishaan, how did you get that?

OpenStudy (anonymous):

but i think my way is more direct

OpenStudy (anonymous):

@ByteMe , follow @myko 's method to get @Ishaan94 's equation

OpenStudy (anonymous):

@myko you have made a mistake

OpenStudy (anonymous):

log_a(b)=a/b

OpenStudy (anonymous):

log_x(y)+log_y(x)=y/x+x/y=

OpenStudy (anonymous):

@myko \[{{\log_x (y)} \over {\log_x (z)} }\neq \log_x{{y} \over {z}}\]

OpenStudy (anonymous):

log_x(y)=only y/x

OpenStudy (anonymous):

now simplify :P

OpenStudy (dumbcow):

first of all you need to understand the change of base formula \[\large \log_{x}y = \frac{\log y}{\log x} = \frac{1}{\frac{\log x}{\log y}} = \frac{1}{\log_{y}x} \] Let a = log_y (x) \[\rightarrow a + \frac{1}{a} = \frac{10}{3}\] \[\rightarrow a^{2} +1 = \frac{10}{3}a\] (ishaans equation)

OpenStudy (anonymous):

y/x+x/y=10/3 =>y^2+x^2/xy=10/3

OpenStudy (anonymous):

yo whatever :P

OpenStudy (anonymous):

@dumbcow , very good explanation . @myko , Now you are right :)

OpenStudy (anonymous):

ya, you right...:) log_y (x) = log_x (x)/log_x(y) so: log_x (x)/log_x(y) +log_x(y) = 10/3 1/log_x(y) +log_x(y)=10/3 log_x(y)^2 +1=10/3log_x(y)^2

OpenStudy (anonymous):

So i'm gettng a=14/3 and a=2 from completing the square. How do i get (x+y)/2 ?

OpenStudy (anonymous):

@ByteMe , check again. I am getting \[a=3\] and \[a={1 \over 3}\]

OpenStudy (dumbcow):

i agree, this implies there are 2 answers for (x+y)/2

OpenStudy (anonymous):

ok ill check. theres only one answer though. Do I put it up?

OpenStudy (anonymous):

Case 1. a= 3 a=logx(y)=3 y=x^(3) Now use x*y = 144 144/x=x^3 144=x^4 x= ±√12 Now substitute this in x*y=144 to get your y Case 2, a= 1/3 Proceed above as in case 1

OpenStudy (anonymous):

nvmd. i see where i made my mistake.

OpenStudy (dumbcow):

i stand corrected...either case yields equivalent x,y values

OpenStudy (anonymous):

@dumbcow ,yes so final answer I am getting \[{13\sqrt{12} }\over 2\]

OpenStudy (anonymous):

On simplifying it becomes or \[13\sqrt(3)\]

OpenStudy (anonymous):

yes!!! thank you very much guys!!!

OpenStudy (anonymous):

@dumbcow , I am not certain about the sign of x and y.

OpenStudy (anonymous):

i'm sorry i have only 1 medal to give :(

OpenStudy (anonymous):

inee meenie mynee moe!!!

OpenStudy (dumbcow):

both x,y are positive

OpenStudy (anonymous):

Yes ofcourse :P It is log :)

OpenStudy (anonymous):

I meant base of log and the quantity inside log has to be positive :)

OpenStudy (dumbcow):

haha :)

OpenStudy (anonymous):

you guys are the best!

OpenStudy (anonymous):

Nope we are just learning while sharing :)

OpenStudy (anonymous):

Nope we are just learning while sharing :)

OpenStudy (anonymous):

Nope we are just learning while sharing :)